# ProSAM: Enhancing the Robustness of SAM-based Visual Reference Segmentation with Probabilistic Prompts

### v2

@@ 7.4 Mathematical Proofs
$f(z+\epsilon)=f(z)+\nabla f(z)^{T}\epsilon+\frac{1}{2}\epsilon^{T}H_{f}(z)\epsilon+R(\epsilon),$
where
@@
$\mathbb{E}_{\epsilon}\Big{[}f(z+\epsilon)\Big{]}=f(z)+\frac{1}{2}\mathbb{E}_{\epsilon}\Big{[}\epsilon^{T}H_{f}(z)\epsilon\Big{]}+O(\sigma^{3}).$
@@
Given that $\mathbb{E}_{\epsilon}[\epsilon_{i}\,\epsilon_{j}]=\sigma^{2}$ if $i=j$ and 00 otherwise, it follows that:
@@ F-0 F-1 F-2 F-3
9.3 Inference Strategies
As mentioned in Section 4.2, the predicted mean prompt $\bm{\hat{\mu}}_{\bm{z}}$ is employed to generate the predicted mask during inference, since the robustness of $\bm{\hat{\mu}}_{\bm{z}}$ can be guaranteed due to the existence of margin between $\bm{\hat{\mu}}_{\bm{z}}$ and the boundary of target prompt region $\mathcal{R}_{I_{r},M_{r},I_{t}}$. However, it is still interesting to see the effectiveness of utilizing randomly sampled prompts from the learned prompt distribution $\bm{z}\sim q_{\phi}(\bm{z}|I_{r},M_{r},I_{t})$ during inference. Therefore, we experimented with 5 different inference strategies, in which we sample $K$ prompts and merge their corresponding $K$ masks in 5 different ways. Specifically, they merge either the logit masks or the binary masks given a threshold of 0.5. From Table 8, all the inference strategies with different ways of merging $K$ masks achieve comparable results to using only the mask prompted by the $\bm{\hat{\mu}}_{\bm{z}}$. This indicates that our learned multivariate prompt distribution does span over the target prompt region $\mathcal{R}_{I_{r},M_{r},I_{t}}$.
Methods Inference Strategies mIOU
ProSAM mean prompt only 75.26
ProSAM max of $K$ logit masks 74.69
mean of $K$ logit masks 75.16
max of $K$ binary masks 74.67
mean of $K$ binary masks 74.92
majority vote of $K$ binary masks 75.05

### v3

@@ 7.4 Mathematical Proofs
$f(z+\epsilon)=f(z)+\nabla f(z)^{T}\epsilon+\frac{1}{2}\epsilon^{T}H_{f}(z)\epsilon+R(\epsilon)$ (13)
, where
@@
$\mathbb{E}_{\epsilon}\Big{[}f(z+\epsilon)\Big{]}=f(z)+\frac{1}{2}\mathbb{E}_{\epsilon}\Big{[}\epsilon^{T}H_{f}(z)\epsilon\Big{]}+O(\sigma^{3}).$ (14)
@@
Given that $\mathbb{E}_{\epsilon}[\epsilon_{i}\,\epsilon_{j}]=\sigma^{2}$ if $i=j$ and 0 otherwise, it follows that:
@@
[…]
Following Equation 14, we observed that the second‐order Taylor terms of $\mathbb{E}[L(z+\epsilon)]$ are identical for any zero‐mean noise with covariance $\sigma^{2}I$. Third‐order terms also vanish because the noise distributions are symmetric and have zero odd moments. The distinction appears first in the fourth‐order term, which depends on the fourth central moment
$m_{4}=\mathbb{E}[\epsilon_{i}^{4}].$
For a Gaussian $\mathcal{N}(0,\sigma^{2})$, one has
$m_{4}^{\mathcal{N}}=3\,\sigma^{4},$
[…]
$m_{4}^{t}=\frac{3\,\nu}{\nu-4}\,\sigma^{4}>3\,\sigma^{4}.$
Recalling that the fourth‐order correction in the expected loss is
$\frac{1}{24}\sum_{i,j,k,\ell}\frac{\partial^{4}L}{\partial z_{i}\partial z_{j}\partial z_{k}\partial z_{\ell}}(z)\,\mathbb{E}[\epsilon_{i}\epsilon_{j}\epsilon_{k}\epsilon_{\ell}],$
[…]
$\frac{m_{4}}{24}\sum_{i}Q_{iiii}+\;\frac{3\,\sigma^{4}}{24}\sum_{i\neq j}Q_{iijj},$
[…]
@@ 8.2 Prompt Distribution Prediction
9 Additional Verification Study
In addition to the studies presented in Section 5.3, we also designed a verification study that does not require training a deep learning model, allowing for a direct comparison between the underlying principles of the variational and non-variational prompt encoders.
For the experiment of learning prompt embedding, we ran 200 experiments to generate 200 prompt embeddings. The initial prompt embeddings are randomly drawn from the normal distribution with a mean of 0 and a standard deviation of 25, and the loss function is formulated as the same loss function as VRP-SAM (see Equation 4). For the experiment of learning variational prompt distribution, we learn a single multivariate prompt distribution following the same formulation and reparameterization trick in Equation 5 and Equation 7 via gradient descent, given the loss function presented in Equation 8. For both experiments, the gradient descent optimization process is stopped when the loss value does not improve by more than 0.0003 over 100 consecutive epochs. For other experimental settings not mentioned above, we follow the same practice as in Section 5.1.